Your firm will pay a dividend of $5.00 per share in perpetuity. The shareholders in your firm have a dividend tax rate of 38 percent. The tax rate on capital gains is 11 percent. The required rate of return on the company's stock is 10.1 percent compounded annually. Your firm has announced that it will no longer pay a dividend but will use the cash to repurchase shares. By how much will stock price change?
Answer #1

Value of stock after repurchase will be = \large \frac{\frac{Dividend }{1+ Dividend Tax}}{1+Capital Gain Tax} * (1+Required Return)

= \large \frac{\frac{5}{1+ 0.38}}{1+0.11} * (1+0.101)

= 3.2641336989 * 1.101

= 3.5938 OR 3.59

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.